Calculate the modulus and the principle value of the argument of the following complex no -3+j4

1
Expert's answer
2022-06-09T11:44:35-0400
"|-3+j4|=\sqrt{(-3)^2+(4)^2}=5"

"\tan \theta=\dfrac{4}{-3}, -\pi<\theta\le \pi"

Quadrant II

"\theta=-\tan^{-1}(\dfrac{4}{3})+\pi"

"|z|=5, Arg(z)=\pi-\tan^{-1}(\dfrac{4}{3})"